BODY FLUIDS CASE STUDY

A 60 year-old women visits her physician with abdominal swelling. It has been determined she has ascites. Paracentesis is performed and patient's blood is drawn after which the specimens are submitted to the lab for testing.

Blood Chemistry Testing

Parameters

Result

Range

Total Protein

6.5 g/dL

(6.0 – 8.5)

LDH

300 U/L

(275 - 630)

Glucose (fasting)

85 mg/dL

(70 - 100)

AST

20 U/L

(10 - 40)

ALT

32 U/L

(10 - 60)

GGT

15 U/L

(0 - 45)

ALP

60 U/L

(45 - 145)

Total Bilirubin

0.6 mg/dL

(0.2 – 1.2)

Peritoneal Fluid Testing

     Examination

Parameters

Result

     Physical

Color

yellow

Clarity

clear

Clots

none observed

     Chemical

Total Protein

2.9 g/dL

LDH

125 U/L

Glucose

70 mg/dL

     Microscopic

Leukocyte count

1200 cells/mL

Cytology exam

no malignant cells observed

Gram stain

no organisms observed

(1) Calculate the fluid-to-serum total protein and fluid-to-serum LDH ratios.

(2) Classify this effusion as a transudate or an exudate.

(3) State at least two physiologic mechanisms which might cause this type of effusion.
Answer #1

Fluid accumulation in the abdominal cavity is called ascites.

Paracentesis is the removal of fluid from the peritoneal cavity through a small surgical incision or puncture made through the abdominal wall under sterile conditions.

(1) Calculate the fluid-to-serum total protein and fluid-to-serum LDH ratios.

Peritoneal fluid total protein = 2.9 g/dL

Serum total protein= 6.5 g/dL

Fluid to serum total protein = 2.9/6.5= 0.45 g/dl

Peritoneal fluid LDH = 125U/L

Serum LDH= 300U/L

Fluid to serum LDH= 125/300=0.42 U/L

-----------------------------------------------------------------------------------------------------

(2) Classify this effusion as a transudate or an exudate.

Transudate is accumulation of fluid due to varied pressure within blood vessels and volume of protein in blood .

Exudate is abnormal accumulation of fluid in abdomen due to injury or any inflammation.

Fluid to serum total protein = 2.9/6.5= 0.45 g/dl (If this is >0.5 then this may be exudate)

Fluid to serum LDH= 125/300=0.42 U/L ( If this is > 0.6 then this may be exudate)

But as these both donot satisfy the criteria for being exudate , this effusion is transudate.

Also transudate appears as clear, pale yellow fluid, the total protein will be < 3 g/dL .Fluid / serum protein will be <0.5 g/dl , fluid / serum LDH will be <0.6 U/L and glucose levels stays same as in blood in transudates.

--------------------------------------------------------------------------------------------------------

(3) State at least two physiologic mechanisms which might cause this type of effusion

Congestive cardiac failure, Nephrotic syndrome and hepatic cirrhosis leads to transudate .